### v1

Abstract
Generative machine learning models have begun to transform protein engineering, yet no principled framework for conditioning on auxiliary information in a plug-and-play manner exists; one may want to iteratively incorporate experimental feedback, or make use of an existing classifier—such as for predicting enzyme commission number—in order to guide the sampling of the generative model to generate sequences with desired properties. Herein, we present ProteinGuide, a rigorous and general framework to achieve just that: through unifying a broad class of protein generative models that includes masked language, (order-agnostic) autoregressive, diffusion and flow-matching models, we provide an approach to statistically condition pre-trained protein generative models. We demonstrate applicability of our approach by guiding each of two commonly used protein generative models, ProteinMPNN and ESM3, to generate amino acid and structure token sequences conditioned on several user-specified properties, namely, enhanced stability and CATH-labeled fold generation.
Generative models are beginning to better enable protein engineers to create bespoke proteins for therapeutic, agricultural and environmental uses [1, 2, 3]. The upshot of any AI-based protein engineering approach is to ultimately specify a protein sequence, or a set of such sequences, even when designing by way of structure. For example, one might first use a backbone structure generative model such as RFdiffusion [2], Chroma [3], or ESM3 [4], followed by use of a backbone-conditioned sequence generative model (also called an inverse-folding model), such as ProteinMPNN [1] or ESM-IF1 [5], to generate sequences likely to fold into the specified backbones. Alternatively, one might forgo generating a backbone structure intermediary and instead generate sequences directly from an “evolutionary” sequence-only model such as ESM3 [4], ProGen [6], EvoDiff [7], DPLM [8] or many others [9, 10]. In all cases, the final step employs a generative model of sequences.
A key ingredient missing from the current protein engineering AI menu is the plug-and-play ability to post hoc “guide” these sequence generative models toward properties of interest, because the technical underpinnings to do so have remained elusive. By guide, we mean, formally, to condition the probability distribution encoded by an existing generative model, $p(x)$, on newly-desired properties of interest, $p(y\in Y|x)$, so as to be able to generate (i.e., sample) desired sequences $x\sim p(x|y\in Y)$. Importantly, we want to do this conditioning in a plug-and-play manner, namely, without re-training the existing model. Although it is currently possible to re-train generative models to achieve property conditioning [11, 12], the plug-and-play characteristics of guidance would enable everyday users to readily adapt existing large, expensive-to-train models of interest to suit their needs, on-the-fly [13, 14, 15].
Figure 1: Equivalence between classes of generative models enables guidance with ProteinGuide. a. Schematic showing how masked language models (e.g., ESM3) and order-agnostic autoregressive models (e.g., ProteinMPNN) can be viewed through the lens of masking diffusion/flow models. This re-interpretation enables the application of predictor-guidance to obtain ProteinGuide. Masking of one sequence position is denoted with the symbol ”?”, and colored dots help to visually tag individual sequences. The sequence being masked during training is M K R S V T; $p_{\theta}(\mathbf{x})$ denotes the pre-trained generative model. The shared training objective involves randomly masking out parts of the sequence and then predicting the masked sequence given the remaining unmasked context b. Contrasting unguided generation with predictor-guidance generation by way of ProteinGuide. $\mathbf{x}$ denotes a sequence and $p_{\theta}(\mathbf{x})$ denotes the pre-trained generative model; $y$ denotes a property of interest that we seek to guide toward using predictive model $p_{\phi}(y\mid\mathbf{x})$. $\Delta t$ denotes change in time in the diffusion process. When guiding, we blend information from the predictive model with that of the pre-trained model in a statistically correct manner. Red letters indicate mutations that are introduced when guiding with the predictive model.
For example, one may want to guide an inverse-folding model to produce proteins with greater or lesser stability than the model naturally generates, or to produce proteins more likely to catalyze a particular reaction, and so forth. Perhaps more importantly, one might want to guide a model using newly acquired experimental measurements, such as catalytic activity for a desired reaction—that is, integrate experimental data in a statistically coherent manner with an existing generative model, without retraining the existing model. Such an ability would enable training-efficient iterative cycles of design, build, test, and learn—a process where experimental feedback guides subsequent AI-based design rounds.
We recently showed for the first time how to correctly and tractably unlock guidance for the two specific classes of generative models of diffusion, and flow matching, on discrete state-spaces [16]. Discrete state-spaces include not only sequences of amino acid, DNA and RNA, but also SMILES strings, graphs, language tokens, structure tokens, or any object represented by a sequence of discrete tokens. Herein, we i) newly show how to use guidance on other commonly-used classes of generative models for discrete spaces, including a large class of Masked Language Models (MLMs) such as ESM3, and order-agnostic auto-regressive models (OA-AR) such as ProteinMPNN, and ii) illustrate in silico application of our method for representative use cases in protein engineering including application to ESM3 and ProteinMPNN. We call our method ProteinGuide.
The essential technical formalism required to derive ProteinGuide, namely to generalize guidance to MLM and OA-AR models, is the connection between how these and diffusion models on discrete state-spaces are trained. Despite that conceptually these models seem to be quite different, it turns out that the loss functions used for training OA-AR models, a broad class of MLMs, as well as masking diffusion and flow models are equivalent (Fig. 1a and Methods). Consequently, herein, we can view a trained MLM or OA-AR model as a diffusion model on a discrete state-space. Correspondingly, we can then leverage our guidance methodology [16] for these models (Fig. 1b and Methods). Although connections between the training objectives of these model classes has been previously noted [17, 18, 19, 20, 21], the practical utility of this connection has been unclear. Here, we show that this connection enables a general framework for guiding a broad class of models. We also note that for autoregressive (AR) models, one can use FUDGE [22] for guidance, which can be viewed as a special case of ProteinGuide for the case of AR models.
To illustrate the potential of ProteinGuide, we apply it to two representative, well-known protein generative models, ProteinMPNN and ESM3. For ProteinMPNN, we trained a regression model on experimental stability measurements [23] with which to guide ProteinMPNN to generate amino acid sequences that encode proteins that are more stable than what ProteinMPNN would do on its own. By employing guidance, we do so in a plug-and-play manner, rather than by retraining the inverse folding model. In our second example, we use ESM3 to generate structure tokens, so as to demonstrate the generality of ProteinGuide to not just amino acid sequences but also other sequences, such as of structure tokens. Specifically, we guide ESM3 to generate backbone structures with specified CATH fold class labels [24], by way of a classifier we trained on CATH data. Although ESM3 was trained to condition on free-form function word tokens, ProteinGuide offers a complementary, plug-and-play approach to conditioning on any property for which we have a predictive model (classifier or regressor), and so is not limited to conditioning on representations or data chosen at the time of ESM3 training.
Importantly, because our approach implements correct statistical conditioning, it preserves the knowledge embedded in the already-trained generative model, while guiding the generation process toward regions of sequence space with desired functional specification. This lies in contrast to heuristic methods such as continuing to train an existing model with new data (i.e., supervised fine-tuning), for which the statistical outcome is not well-defined, and for which important information from the already-trained model may be forgotten. Moreover, we do so without having to adjust the pre-existing model’s parameters.
Direct Preference Optimization (DPO [26, 11]) and related work enables updating of pre-trained models, although up until recently, these methods have been applicable only to models with tractable likelihoods, thereby restricting them largely to autoregressive models. Wallace et al. [27] extended DPO to diffusion models on continuous state-spaces by way of approximation, while Borso et al. [28] explored how to similarly do so for these models on discrete state-spaces. Importantly, the connections previously mentioned between MLMs, diffusion and flow matching models on discrete state-spaces, and order-agnostic autoregressive models, means that we can also generalize the aforementioned approximate DPO to these same model classes. Herein, we focus our study on plug-and-play guidance, rather than DPO-style approaches, although the latter could prove equally useful.
[…]
On the whole, we observed that guidance, as expected, led to desired conditioning on each of our tasks. Also as expected, when guiding ESM3 with CATH labels, more fine-grained levels and also rarer classes were more challenging—that is, had lower success rates (Fig. 2b). We expect the quality of generated structures would improve concomitantly with improved backbone generative models, independent of guidance.
In conclusion, we introduced ProteinGuide, a broadly applicable framework that enables conditioning of pre-trained protein generative models on any property of interest that can be encoded in a predictive model, without having to re-train the pre-trained model. An important application enabled by our method is to incorporate feedback from experimental measurements into the generative process.
Figure 2: In silico demonstration of ProteinGuide with ProteinMPNN and ESM3. a. Generating structure-conditioned protein sequences from ProteinMPNN with no guidance (teal) and with ProteinGuide for stability (pink), across eight proteins (horizontal axis). The height of each bar represents the percentage of sequences predicted to be at least as stable than wildtype ($\Delta\Delta G\geq 0$), while the hatched sub-portion indicates the subset of those stable sequences that also correctly fold into the desired structure (RMSD $\leq 2$Å). b. Generating backbone structure tokens with ESM3 with no guidance (teal), and with fold class ProteinGuided ESM3 (pink), across different CATH hierarchy levels. The height of each bar shows the percentage of structures correctly reclassified to their target fold class, while the hatched sub-portion indicates the subset that are also designable (self-consistency RMSD $\leq 2$Å). Results are grouped by CATH level. For level C, there are three classes: mainly $\alpha$, mainly $\beta$, and mixed $\alpha/\beta$. For level A and T, we grouped the classes for each level by their frequency of occurrence in the AFDB (Common, Regular, Rare) and calculated the metrics within each group following Geffner et al. [29]. c. Examples of randomly chosen proteins generated by guiding ESM3 with three example CATH classes.
Acknowledgments
This work was funded in part by the the U.S. Department of Energy, Office of Science, Office of Biological and Environmental Research, Lawrence Livermore National Laboratory BioSecure SFA within the Secure Biosystems Design program (SCW1710), and also the Office of Naval Research (ONR) under grant N00014-23-1-2587.
Methods
Background on discrete state-space diffusion and flow models
[…]
$p(x_{t+dt}=\tilde{x}|x_{t}=x)=\delta_{x,\tilde{x}}+R_{t}(x,\tilde{x})dt+O(dt^{1+\epsilon}).$ (1)
As an example, Campbell et al. [32] defines a forward noising process that gradually corrupts data by transitioning to a noise distribution. The model is trained to reverse this process, learning rates that can denoise corrupted samples.
There are also a number of works that extend flow matching, which were originally developed for continuous state-spaces [35, 36, 37], to discrete state-spaces [38, 39], also leveraging CTMCs, but differ from diffusion models in how the CTMCs are defined, which also induce different training objectives. These models specify a conditional flow that defines how probability mass should move from a noise distribution to the data distribution. The model learns to generate samples by following the flow trajectories. Campbell et al. [38] show that the rates used in generative sampling can be obtained by training a neural network with parameters $\theta$, $p^{\theta}(x_{1}|x_{t})$, to approximate the true denoising distribution using the standard cross-entropy loss,
$\mathcal{L}_{\text{FM}}=\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),\,t\sim p(t),\,x_{t}\sim p_{t|1}(x_{t}|x_{1})}\left[\log p^{\theta}(x_{1}|x_{t},t)\right],$ (2)
where $x_{1}$ is the clean data, $p(t)$, referred to as the noise schedule, is a distribution with full support on the interval $[0,1]$, and $x_{t}$ is the noised data sample from the forward noising process $p_{t|1}(x_{t}|s_{1})$. [^1] For all models, we follow the convention of flow-based models, where $t=1$ is the target (training) distribution, and $t=0$ is the noise distribution. As shown in Campbell et al. [38], the cross entropy loss can be understood as a simplification to the ELBO used to train continuous-time discrete diffusion models [32]. Given such a trained denoising model $p^{\theta}(x_{1}|x_{t})$, the unconditional rate matrix used in sampling is defined as
$R^{\theta}_{t}(x_{t},x_{t+dt})=\mathbb{E}_{x_{1}\sim p^{\theta}(x_{1}|x_{t})}\Big{[}\overrightarrow{R_{t}}(x_{t+dt},x_{t}|x_{1})\Big{]},$ (3)
where $\overrightarrow{R_{t}}(x_{t+dt},x_{t}|x_{1})$ is the forward noising rate matrix induced by $p(x_{t}|x_{1})$.
A popular instantiation of both discrete state-spaces diffusion and flow models uses masking as the noising process, where tokens in a sequence are progressively replaced with mask tokens during the forward process and the model learns to recover the original tokens during the reverse process. For example, the masking conditional flow can be defined as $p^{\text{mask}}(x_{t}|x_{1})=\text{Cat}(t\delta\{x_{1},x_{t}\}+(1-t)\delta\{M,x_{t}\})$. This formulation has connections to masked language models and order-agnostic autoregressive models, which we will describe in detail next.
Unifying connections between generative models on discrete state-spaces
Connections among masked discrete diffusion models, masked language models (MLM) and order-agnostic autoregressive models (OA-AR) were initially pointed out by Austin et al. [17] and Hoogeboom et al. [18]. Recent works [19, 20, 21] further clarified these connections, including the connections to CTMC-based formulation of diffusion and flow models [32, 34, 38], and showed that the continuous-time ELBO used to train masked diffusion models is essentially a weighted sum of cross entropy losses predicting the masked states from unmasked positions, where the proportion of masked positions ranges from 0 (complete masking) to 1 (no masking). For particular choices of the noise schedules, these training objectives can be shown to be equivalent to the training objective of order-agnostic autoregressive models [40], (generative) masked language models [17], or discrete flow matching models [38, 39] under a masking process. Since ESM3 [4] and ProteinMPNN [1] were trained under the MLM and OA-ARM objectives respectively, we can view them as discrete state-space diffusion and flow models, and sample from them as such, leveraging the generative rate matrices from Equation 3 induced by the trained denoising model, $p^{\theta}(x_{1}|x_{t})$. This reformulation not only provides a principled way to sample from these models, but also offers additional sampling flexibility, including, but not limited to, the usage of different noise schedules from training and corrector-sampling [38, 39], which have been shown to sometimes improve sampling results compared to prior heuristics [41] for sampling from generative masked language models [39]. In addition, this reformulation unlocks the capacity to condition the generative process via guidance, which we introduce in Section Background on diffusion guidance.
Here, we provide a self-contained derivation of the equivalence between the loss functions of discrete flow matching, masked language models, and order-agnostic autoregressive models.
Equivalence between flow matching on discrete state-space and masked language modeling
Recall from Equation 2 that the flow matching loss is
$\mathcal{L}_{\text{FM}}=\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),\,t\sim p(t),\,x_{t}\sim p_{t|1}(x_{t}|x_{1})}\left[\log p^{\theta}(x_{1}|x_{t},t)\right],$ (4)
where $x_{1}$ refers to clean (un-noised) data, $p(t)$ is the noise schedule for $t\in[0,1]$, and $x_{t}$ the noised data. Campbell et al. [38] show that when the forward noising process is applied independently across each dimension of $x_{1}$, the denoising model, $p^{\theta}(x_{1}|x_{t},t)$, can be learned as a factorized over each dimension, $i\in\{1,2,...,D\}$, of the data:
$\mathcal{L}_{\text{FM}}$ $=\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),t\sim p(t),,x_{t}\sim p_{t|1}(x_{t}|x_{1})}\left[\log\prod_{i=1}^{D}p^{\theta}(x_{1}^{i}|x_{t},t)\right].$ (5)
Our goal is to demonstrate that this loss function is equivalent to the loss function used by a class of masked language models (MLMs), that we will refer to as “generative masked language models”. The key distinction between generative masked langauge models and their non-generative counterparts, is that during training, generative MLMs mask tokens with a masking probability, $t\in[0,1]$, sampled from a distribution $p(t)$ that has full support on the interval $[0,1]$. This requirement is met by MLMs such as ESM3, but, importantly, is not met by BERT-style MLMs, such as ESM2 as these models use a fixed percentage of tokens that are masked at each step. Letting, $x_{t}$ refer to the masked sequence, we have that the generative masked language model objective function is
$\mathcal{L}_{\text{MLM}}$ $=\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),t\sim p(t),,x_{t}\sim p_{t|1}(x_{t}|x_{1})}\left[\log\prod_{i=1}^{D}p^{\theta}(x_{1}^{i}|x_{t})\right],$ (6)
where we can see that the key difference between Equation 6 and Equation 2 is that the latter has a denoising neural network with a dependence on $t$. We will now show that when a masking noise process is used for flow matching, the denoiser’s dependency on $t$ can be dropped, making the two loss functions equivalent.
Let $x_{1}=(x_{1}^{1},x_{1}^{2},...,x_{1}^{D})$ be our original sequence from the data distribution. For a masking noise process we define binary masking variables $m_{t}^{i}\in\{0,1\}$ for each position $i$ at time $t$, where $m_{t}^{i}=1$ indicates token $i$ is masked at time $t$ and $m_{t}^{i}=0$ indicates token $i$ is not masked at time $t$. Using ‘?’ to denote a masked state, the state $x_{t}$ is then defined as:
$x_{t}^{i}=\begin{cases}x_{1}^{i}&\text{if }m_{t}^{i}=0\\ \text{?}&\text{if }m_{t}^{i}=1.\end{cases}$ (7)
Let us additionally define:
  1. 1.
    $M_{t}=\{i:x_{t}^{i}=\text{?}\}$ as the set of masked positions at time $t$.
  2. 2.
    $U_{t}=\{i:x_{t}^{i}\neq\text{?}\}$ as the set of unmasked positions.
For any position $i\in U_{t}$, we know with certainty that $x_{1}^{i}=x_{t}^{i}$ by the definition of our forward process. The key observation is that the time variable $t$ only influences which tokens are masked (the masking pattern), but once we observe $x_{t}$, we have complete information about which positions are masked and which are not. The time variable, therefore, provides no additional information needed to predict the clean data. Using the mathematical argument from Gat et al. [39], we can prove this as follows:
$p(x_{1}\mid x_{t},t)$ $=\frac{p(x_{t},t\mid x_{1})p(x_{1})}{\sum_{\tilde{x}_{1}}p(x_{t},t\mid\tilde{x}_{1})p(\tilde{x}_{1})}$ (8)
$=\frac{\cancel{\big{[}\prod_{i\in M_{t}}(1-t)\prod_{i\in U_{t}}t\big{]}}\big{[}\prod_{i\in U_{t}}\mathbb{I}(x_{t}^{i}=x_{1}^{i})\big{]}p(t)p(x_{1})}{\sum_{\tilde{x}_{1}}\cancel{\big{[}\prod_{i\in M_{t}}(1-t)\prod_{i\in U_{t}}t\big{]}}\big{[}\prod_{i\in U_{t}}\mathbb{I}(x_{t}^{i}=\tilde{x}_{1}^{i})\big{]}p(t)p(\tilde{x}_{1})}$ (9)
$=\frac{\big{[}\prod_{i\in U_{t}}\mathbb{I}(x_{t}^{i}=x_{1}^{i})\big{]}p(x_{1})}{\sum_{\tilde{x}_{1}}\big{[}\prod_{i\in U_{t}}\mathbb{I}(x_{t}^{i}=\tilde{x}_{1}^{i})\big{]}p(\tilde{x}_{1})}$ (10)
$=p(x_{1}\mid x_{t}),$ (11)
where $\mathbb{I}$ is the indicator function. Thus, we can rewrite the masking flow matching loss function as:
$\mathcal{L}_{\text{FM}}$ $=\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),\,t\sim p(t),\,x_{t}\sim p_{t|1}(x_{t}|x_{1})}\left[\log\prod_{i=1}^{D}p^{\theta}(x_{1}^{i}|x_{t},t)\right]$ (12)
$=\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),\,t\sim p(t),\,x_{t}\sim p_{t|1}(x_{t}|x_{1})}\left[\log\prod_{i=1}^{D}p^{\theta}(x_{1}^{i}|x_{t})\right]$ (13)
$=\mathcal{L}_{\text{MLM}}.$ (14)
Equivalence between flow matching on discrete state-space and order-agnostic autoregressive models
Order-agnostic autoregressive models (OA-AR) differ from standard autoregressive models by training on arbitrary permutations of the sequence ordering. We will now demonstrate that under specific conditions, the discrete flow matching objective is equivalent to the order-agnostic autoregressive objective.
Let us consider a sequence $x_{1}=(x_{1}^{1},x_{1}^{2},...,x_{1}^{D})$ from our data distribution. In an order-agnostic autoregressive model, we sample a permutation $\sigma$ of indices $\{1,2,...,D\}$ and factorize the joint distribution as:
$p^{\theta}(x_{1})=\prod_{i=1}^{D}p^{\theta}(x_{1}^{\sigma(i)}|x_{1}^{\sigma(1)},x_{1}^{\sigma(2)},...,x_{1}^{\sigma(i-1)})$ (15)
The OA-AR objective is to maximize the expected log-likelihood over all possible permutations:
$\mathcal{L}_{\text{OA-AR}}$ $=\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),\sigma\sim\text{Unif}(\text{Perm}(D))}\left[\sum_{i=1}^{D}\log p^{\theta}(x_{1}^{\sigma(i)}|x_{1}^{\sigma(1)},x_{1}^{\sigma(2)},...,x_{1}^{\sigma(i-1)})\right]$ (16)
$=D\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),\sigma\sim\text{Unif}(\text{Perm}(D)),i\sim\text{Unif}(\{1,2,\ldots,D\})}\left[\log p^{\theta}(x_{1}^{\sigma(i)}|x_{1}^{\sigma(1)},x_{1}^{\sigma(2)},...,x_{1}^{\sigma(i-1)})\right]$ (17)
$=D\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),i\sim\text{Unif}(\{1,2,\ldots,D\}),\sigma\sim\text{Unif}(\text{Perm}(D))}\left[\log p^{\theta}(x_{1}^{\sigma(i)}|x_{1}^{\sigma(1)},x_{1}^{\sigma(2)},...,x_{1}^{\sigma(i-1)})\right]$ (18)
$=D\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),t\sim\text{Unif}([0,1]),x_{t}\sim p_{t|1}(x_{t}|x_{1})}\left[\log p^{\theta}(x_{1}\mid x_{t})\right]$ (19)
$=D\mathcal{L}_{\text{FM}},$ (20)
where we go from Equation 18 to Equation 19 by recognizing that randomly picking the number of positions to mask and sampling a random permutation is equivalent to picking a random time and sampling the forward noising process for that time. Thus, we have shown that the order-agnostic autoregressive modeling loss is equivalent, up to a constant, to a masking discrete flow matching loss with a uniform noise schedule.
Relationship to autoregressive models and discrete time discrete state-space diffusion models
Autoregressive (AR) models represent another important class of generative models for discrete spaces that can be connected to the diffusion framework. AR models factorize the joint distribution sequentially as $p(x_{1})=\prod_{i=1}^{D}p(x_{1}^{i}|x_{1}^{1},x_{1}^{2},...,x_{1}^{i-1})$ , generating tokens one at a time in a fixed order. Austin et al. [17] showed that AR models can be interpreted as discrete space diffusion models with a particular form of noising process.
For AR models, Yang and Klein [22] introduced FUDGE, a method that performs guidance by modulating each conditional distribution using a learned classifier evaluated on the partially completed sequence. FUDGE can be viewed as a special case of ProteinGuidewhen applied specifically to AR models. While discrete time diffusion models are trained with a different objective than their continuous-time counterparts, they can still benefit from our guidance methodology. The learned denoising model from a discrete-time training objective can be used to induce a rate matrix for sampling by assuming smooth interpolation between discrete time points. However, both theoretical analysis and empirical results [32, 34, 38, 39] suggest that models trained with continuous-time objectives often achieve superior performance in practice.
This relationship between different model classes highlights the generality of ProteinGuide as a framework that can accommodate various generative architectures, including AR models, discrete-time diffusion models, and the continuous-time formulations we primarily focus on in this work.
Background on diffusion guidance
Guidance is a useful technique that enables conditional generation with diffusion and flow models. Originally developed for continuous state-space models, guidance allows steering the generative process toward samples with desired properties without requiring conditional retraining [13, 15, 14]. In continuous state-space diffusion models, the generation process involves sampling from a reverse stochastic differential equation (SDE) that gradually transforms noise into structured data. To condition this process on a desired property $y$, it has been shown that one can leverage Bayes’ theorem to obtain the conditional score function:
$\nabla_{x_{t}}\log p_{t}(x_{t}|y)=\nabla_{x_{t}}\log p_{t}(x_{t})+\nabla_{x_{t}}\log p_{t}(y|x_{t})$ (21)
This formulation, termed classifier guidance, combines the unconditional score $\nabla_{x_{t}}\log p_{t}(x_{t})$ with the score of a predictor model $\nabla_{x_{t}}\log p_{t}(y|x_{t})$. The strength of guidance can be controlled by introducing a parameter $\gamma$:
$\nabla_{x_{t}}\log p_{t}(x_{t}|y)=\nabla_{x_{t}}\log p_{t}(x_{t})+\gamma\nabla_{x_{t}}\log p_{t}(y|x_{t})$ (22)
This approach provides several advantages: (ii) it enables conditioning an unconditional model without retraining, (ii) the guidance strength $\gamma$ can be adjusted at inference time, and (iii) it allows composition of multiple guidance signals. Ho and Salimans [42] later introduced classifier-free guidance, which achieves similar results without an explicit classifier by combining a conditionally trained model and an unconditional model, which in practice are typically trained jointly. Classifier-free guidance has been shown to improve sample quality [42] compared to using only the conditionally trained model, and recent work on autoguidance showed further improved sample quality by combining a trained model with a more inferior version of itself [43].
While guidance, either classifier-based, classifier-free or autoguidance, has proven extremely valuable for continuous state-space models, extending these concepts to discrete state-spaces poses unique challenges, since the score functions are not defined and these formulations cannot be directly applied. We sidestep the need for a score function by instead returning to Bayes’ theorem, which, in the context of diffusion on discrete state-spaces, dictates that we must effectively compute
$p(x_{t+dt}|x_{t},y)=\frac{p(y|x_{t+dt},x_{t})p(x_{t+dt}|x_{t})}{\sum\limits_{x^{\prime}_{t+dt}}p(y|x^{\prime}_{t+dt},x_{t})p(x^{\prime}_{t+dt}|x_{t})},$ (23)
where $x_{t+dt}$ and $x_{t}$ are separated by a infinitesimal time step $dt$, and $y$ is the desired property we wish to condition on. In a $D$-dimensional discrete state-space where each dimension has cardinality $S$, there are $S^{D}$ terms in the normalizing constant of Equation 23, rendering it generally intractable. Each term arises from the fact that, without constraints, any state can be reached from any other state. Intuitively, tractability can only be achieved if the number of terms is not exponential, thereby implying some constraints on the allowed state transitions. Next, we shall see how the continuous-time formulation of discrete state-space diffusion and flow models enables us to tractably compute the normalizing constant by imposing such constraints, without losing any model expressibility.
Guidance in discrete state-spaces
[…]
To perform predictor guidance in this setting, we derive a principled approach to adjust these rates based on a predictor model $p(y|x_{t})$ that evaluates how likely a state $x_{t}$ is to have the desired property $y$. Specifically, we show that the conditional rates can be obtained as:
[…]
for any state $\tilde{x}\neq x$. This formulation directly implements Bayes’ theorem at the level of transition rates, modulating transitions toward states that have higher likelihood of possessing the desired property.
Similar to continuous state-space guidance, we can control the strength of guidance with a parameter $\gamma$:
[…]
, or derive predictor-free guidance discrete state-spaces:
$R_{t}^{(\gamma)}(x,\tilde{x}|y)=R_{t}(x,\tilde{x}|y)^{\gamma}R_{t}(x,\tilde{x})^{1-\gamma}$ (26)
For computational efficiency, especially with large state-spaces, we introduced Taylor-approximated guidance (TAG), which approximates the likelihood ratio using a first-order Taylor expansion inspired by Grathwohl et al. [44]:
$\log\frac{p_{\phi}(y|\tilde{x},t)}{p_{\phi}(y|x,t)}\approx(\tilde{x}-x)^{T}\nabla_{x}\log p_{\phi}(y|x,t)$ (27)
requiring only one forward and one backward pass of the predictor model instead of $D\times(S-1)+1$ forward passes, making estimation of the guide-adjusted rates $O(1)$ rather than $O(D\times S)$. With multiple particles, TAG can be further improved by acting as a proposal distribution in a Sequential Monte Carlo (SMC) algorithm similar to Wu et al. [45]. While there are other recently proposed approaches for guiding discrete diffusion models [46, 47, 48], ProteinGuide has the benefit of being both exact and general, applicable to both diffusion and flow-based models, making it a general and flexible framework for guiding a broad class of discrete state-space generative models.
Stability-guided inverse folding
We demonstrated how assay-labeled experimental data could be used to condition generative models for protein sequences, improving the success rates for such models in generating sequences with the desired function. To illustrate this use case, we guided ProteinMPNN with a large dataset of experimental folding stability measurements on mini-proteins [23] to generate sequences with improved stability. Stability is reported in real-valued units of $\Delta\Delta G$, which quantifies the difference in stability from the wild-type sequence. A protein is considered stable if it is at least as stable as the wild-type (i.e., $\Delta\Delta G\geq 0$). To generate sequences with improved stability, we followed the experimental setup from Nisonoff et al. [16]. Specifically, we use ProteinMPNN [1] as the inverse folding model, $p(\mathbf{s}\,|\,\mathbf{x})$, where $\mathbf{s}$ corresponds to a protein sequence and $\mathbf{x}$ is the desired backbone structure. Then we apply guidance with the stability predictor from Nisonoff et al. [16], $p(\Delta\Delta G\,|\,\mathbf{s},\mathbf{x})$, to obtain a stability-conditioned inverse folding model, $p(\mathbf{s}\,|\,\mathbf{x},\Delta\Delta G\geq 0)$.
We applied this set-up on the same eight randomly chosen proteins from Nisonoff et al. [16]. For each protein, we sampled 100 sequences using either ProteinMPNN or ProteinMPNN with guidance. For each generated sequence, we define success based on two critera: (1) the generated sequence should fold into the desired structure, $\mathbf{x}$, and (2) the generated sequence should be at least as stable as the wild-type sequence, $\Delta\Delta G\geq 0$. We used AlphaFold 2 [49] to predict the structure of the generated sequences and compare the root mean square deviation (RMSD) between the predicted structure and the desired structure $\mathbf{x}$. We consider a sequence to achieve the desired fold if the RMSD is less than or equal to 22Å following previous evaluation criteria [38]. To evaluate the folding stability of our sequences, we use the $\Delta\Delta G$ regression model $p(\Delta\Delta G\,|\,\mathbf{s},\mathbf{x})$. ProteinMPNN sampling was performed with three different temperatures: $1.0$, $0.1$, and $0.01$ and we report the results for the temperature that achieved the highest success rate. For guided generation, we used a single temperature of 0.1, 200 sampling steps, sampling stochasticity of 20 and guidance factor of 100 for all proteins. On the whole, we observed that across all proteins, performing guidance with ProteinMPNN consistently improves its success rates (Fig. 2a).
Fold class-guided backbone generation
First, we provide an overview of the experimental set-up for guiding ESM3 backbone structure generation with fold class labels from CATH. The purpose of this experiment to illustrate how functional annotations could be used to condition the sampling process of a zero-shot generative model without requiring modifications to the original model. We illustrate this by guiding ESM3 to sample protein backbone structures conditioned on fold class labels from CATH, a hierarchical protein structure classification scheme. This enables both high-level specification of overall secondary structure compositions and more fine-grained descriptions about particular folds. Concretely, we trained a fold classifier, $p(\mathbf{c}\,|\,\mathbf{x})$ that predict the CATH labels, $\mathbf{c}$, from the discrete structure tokens outputted by the structure encoder of ESM3, $\mathbf{x}$, on the set of CATH-labeled protein structures from the Protein Data Bank (PDB). We combined this fold class predictor with the structure generation track of ESM3, $p(\mathbf{x})$, to obtain a fold class conditioned backbone structure generative model, $p(\mathbf{x}\,|\,\mathbf{c})$.
Now we describe additional details of the fold class predictor. Since ESM3 operates on discrete structure tokens, we need a predictor which predicts the fold class label of a protein from its structure token representation. To this end, we first trained a fold class predictor that takes in discrete structure tokens as inputs. The model architecture consists of an encoder-only Transformer trunk. The final per-position embedding is mean-pooled, then fed into three parallel linear layers, each projecting the embedding to a vector with the dimension equal to the number of classes for each of the three levels of CATH hierarchies: C (“Class”), A (“Architecture”) and T (“Topology/Fold”). The C level describes the overall secondary structure-content of a structure (e.g., “mainly $\alpha$”, “mainly $\beta$”, and “mixed $\alpha/\beta$), the A level roughly clusters structures by the overall arrangement of secondary structure in space (e.g.“beta barrel”, “roll”), and the T level assigns structures by how the secondary structure elements are connected and arranged (e.g.“immunoglobulin-like”, “TIM barrel”) [24]. The model was trained on the PDB dataset following the same pre-processing procedure as in Geffner et al. [29] with a total of 214,564 structures, categorized into 5 C classes, 43 A classes, and 1,336 T classes. We fed each structure through the structure encoder of ESM3 to obtain its structure token representation. The dataset is randomly divided into training, validation, and test sets at a ratio of 8:1:1, ensuring that at least one protein from each class is included in the test set whenever possible. The model is trained with a standard cross-entropy loss applied to each of the CATH level. For each protein, the loss is calculated for each level of its available CATH label. We did not systematically tune the hyper-parameters of the model. The trained fold class predictor achieves comparable accuracies on the test set as the predictor from Geffner et al. [29], which takes structure coordinates as inputs as opposed to structure tokens. The fold classifier used for guidance, $p(\mathbf{c}\,|\,\mathbf{x_{t}})$ needs to take in partially masked inputs. We initialized the model weights of this “noisy” classifier with those of the unnoised classifier, and trained it on the same training dataset as the unnoised classifier.
We applied this set-up across the three hierarchies of fold class labels, following the evaluation set-up of Geffner et al. [29]. We sampled backbone structures conditioning on different classes across three levels of the CATH hierarchies. We drew 100 backbone structure samples for each C-level class, 30 samples for each A-level class, and 30 samples for each of the 30 randomly selected T-level classes, using either the unconditional structure generation track of ESM3, $p(\mathbf{x})$, or ESM3 with fold class guidance, $p(\mathbf{x}\,|\,\mathbf{c})$. For each class, for every sample, we sampled its length uniformly among the lengths present for the proteins of that class in the AFDB, with a minimum length of 50 and maximum length of 275. For unconditional generation, we used the default sampling setting of ESM3 with a temperature of 0.7 and $L$ sampling steps, where $L$ is the length of the protein sequence. We also experimented with using a lower sampling temperature and found that it results in ESM3 producing predominantly mainly $\alpha$ proteins. For guided generation, we used a temperature of 0.3, 200 sampling steps, sampling stochasticity of 20 and guidance factor of 100.
We evaluated the generated structures on two criteria: (i) whether the fold classifier, $p(\mathbf{c}\,|\,\mathbf{x})$, predicts the desired class $\mathbf{c}$ as the most likely class for the generated structure $\mathbf{x}$, and (ii) whether the structures are considered designable, where designability is calculated following the procedure in Yim et al. [50]: for each generated backbone, eight sequences are sampled using ProteinMPNN [1] with a temperature of 0.1, then ESMFold [51] is used to predicted the structure of each generated sequence. The RMSD is calculated between the generated structure and the predicted structure, and a generated backbone is considered designable if the lowest RMSD—referred to as the self-consistency RMSD (scRMSD)—among the eight sampled sequences is $\leq 2$Å. For the purpose of computing aggregated metrics across multiple classes, we grouped the classes for each level by their frequency of occurrence in AFDB and calculated the reclassification accuracy and the fraction of designable proteins within each group. The frequency groups are defined in the same way as in Geffner et al. [29].
On the whole, we observed that across all levels and groups, guidance is able to steer the generation towards the desired classes compared to its unconditional counterpart (Fig. 2b). Generally, we observed that the reclassification accuracy is lower for the A and T level and especially so for the rarer classes. Considering that the A and T level fold class labels are more fine-grained and the constraints are harder to satisfy, and increasingly so for the rarer classes given the imbalanced nature of the fold class distribution, this result is expected. We also observed that among the samples that are classified correctly, only a moderate fraction is designable. This aligns with previous observations that ESM3 tends to produce backbones that are less designable, possibly due to the composition of its training set [29]. We also note that increasing designability does not necessarily correlate with a better capability in generating functional backbone structures, as it has been observed that many functional proteins, e.g., those in the CATH databases, are not designable by the standard definition, and backbone generation models with high designability can systematically miss parts of the structure space that are enriched for these functional proteins [52]. Since guided generation depends crucially on the capability of the underlying unconditional model, we can expect structure generation to be improved as better structure generative models are developed.
Additional discussion of related work
Several concurrent works have proposed techniques for steering discrete state-space diffusion models. These approaches can be broadly categorized into guidance-based methods (which modify the sampling process at inference time) and fine-tuning approaches (which update the parameters of the pre-trained model).
Li et al. [46] proposes a derivative-free sampling technique for reward maximization in both continuous and discrete spaces using Sequential Monte Carlo (SMC). Their approach uses the unconditional model as the proposal distribution and reweighs multiple particles with the reward function. Singhal et al. [48] builds upon and generalizes the methods in Wu et al. [45] and Li et al. [46] to enable more flexible sampling hyperparameters. Schiff et al. [47] presented a discrete-time analog of the continuous-time guidance method in Nisonoff et al. [16]. As noted in Schiff et al. [47], in discrete time, the normalizing constant in Equation 23 is still intractable, necessitating additional assumptions to achieve tractability, whereas Nisonoff et al. [16] shows that the continuous-time formulation of discrete diffusion models provides exactly the tractability necessary to perform exact guidance. In addition, the continuous-time formulation naturally enables guidance for discrete state-space flow matching models, whereas the discrete-time formation does not. Additionally, the continuous-time formulation naturally extends to discrete state-space flow matching models. Lee et al. [53] observed that when applying stronger guidance, sampling can be improved with SMC, aligning with challenges in low-temperature sampling noted in Du et al. [54] and Ingraham et al. [3]. For AR models, Zhao et al. [55] proposed an SMC-based steering method that learns intermediate twist functions via contrastive learning, similar to learning time-dependent predictors in diffusion guidance.
Recent work has also explored updating model parameters to incorporate conditioning rather than modifying the sampling process. Wang et al. [56] introduced DRAKES, which uses the Gumbel-Softmax trick to make discrete diffusion trajectories differentiable and enable backpropagation of rewards through entire sampling paths. Concurrently, Rector-Brooks et al. [57] proposed DDPP, which frames the steering problem as sampling from a Bayesian posterior where the pre-trained model serves as the prior and the reward acts as the likelihood. Their approach offers three variants with different estimation strategies. Both fine-tuning approaches establish connections to Nisonoff et al. [16] while offering complementary perspectives on solving the steering problem. For AR models, DPO [26] has emerged as a popular technique for preference-based fine-tuning, which recent work has also extended to discrete space diffusion models [28].